OpenStudy (anonymous):

Factor the trinomial 5s^2-25s+30

OpenStudy (anonymous):

s2-5s+6=0 s2-3s-2s+6=0 s(s-3)-2(s-3)=0 (s-2)(s-3)=0 s=2 or 3

OpenStudy (anonymous):

This is factoring the trinomail, not working for S value